Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR16 in R80 @ 15:30 pm on Tuesday 26th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • Actions from last meeting
    • DS readback current requires calibration (DL)
    • Install FC monitoring in ISIS MCR (Pierrick,Paolo)
    • Install SSH Enable/Disable mimic : Pierrick
  • MOM report
  • Magnet & Channel commissioning
    • Plans for magnet work this week
  • Run plans (dependent on magnet plans)
    • run plan for tonight
  • Shift training
    • Beamline training for Dan Kaplan on Wednesday - late morning / early afternoon
    • Controlled access refresh for Dan on Wednesday - morning
  • DONM : 26th July in CR16

Minutes

Present : SB, AL, CW, VB, TM, YS, AK, MT, DK, DN, JG, HN, DR

  • Actions from last meeting
    • DS readback current requires calibration (DL) : Stands
    • Install FC monitoring in ISIS MCR (Pierrick,Paolo) : Stands
    • Install SSH Enable/Disable mimic : Pierrick : Stands
  • MOM report
    • All straight track data taken
    • FC was brought up and data taken - however not in right direction.
    • FC was taken down, swapped around and brough up again and run successfully
    • No problems with FC operation
      • cold mass temp drifted up over previous week during hot weather. Hall AC units switched on and cold mass temp stabilised.
    • Hysteresis measurements were done and Paolo is analysing them
    • Running smooth except for failure of target computer failure which took 24 hours to recover from
      • Discussion : how many mahcines should we define hot spares for?
    • Henry: we keep having water trips on the magnets. Need to put magnet trips in the alarm handler.
    • Henry: ISIS water leak without alarms.
    • Ken: we need to write down a list of things we need monitored, prioritise and implement.
    • Ken: praise the response of shift crews and experts to the issues over the last couple of weeks.
    • Readiness
  • Magnet & Channel commissioning
    • Plans for magnet work this week
      • Not where we hope to be now. But will be by tomorrow.
      • Planned to ramp SSU, then FC in 3 stages up to a maximum of 50 A stopping at each stage for an hour to monitor. SSU is now up to
        current and the FC has reached the first waiting point. Will see what we get by end of day today. Have arranged things so we have
        magnet expert cover from 8 - midnight. On graveyard, shifters will take data and babysit the magnets.
  • Run plans (dependent on magnet plans)
    • run plan for tonight
      • Take diffuser data plus alignment data
  • Shift training
    • Beamline training for Dan Kaplan on Wednesday - late morning / early afternoon
    • Controlled access refresh for Dan on Wednesday - morning
  • AOB
    • Controlled entry with beamstop not shut has been escalated to an SHE Near Miss.